OpenStudy (anonymous):

A bridge hand is made up of 13 cards from a deck of 52. Find the probability that a hand chosen at random contains at least 3 nines.

OpenStudy (perl):

Let X = number of nines in a 13 bridge hand We want P( X >= 3) . The complement approach would be easier. P( X>= 3 ) = 1 - P( X <3 ) = 1 - P ( X <= 2 ) = 1- ( P(X=0 + P(X=1) + P(X=2) ) $$ \large \rm { P(X \leq 2 ) = 1 - \left ( \frac{ \binom {48}{13} \cdot \binom{4}{0}}{\binom {52}{13}} + \frac{ \binom {48}{12} \cdot \binom{4}{1}}{\binom {52}{13}} + \frac{ \binom {48}{11} \cdot \binom{4}{2}}{\binom {52}{13}} \right ) } $$

OpenStudy (anonymous):

Thank you so much Perl!!!! This is very helpful!!!!!! I have become your fan :)

OpenStudy (perl):

also i might clarify a bit on how i got the fractions, and i am using binomial notation n choose r \( \binom n r \) $$ \Large \rm { \frac{ \binom {not ~nines }{\#} \cdot \binom{nines}{\#}}{\binom {52}{13}} } $$

OpenStudy (perl):

there are 48 'not nines' , and 4 nines